### v1

# The early r-process nucleosynthesis scenarios

Abstract
I compare seven actively studied r-process nucleosynthesis scenarios against observed properties of r-process elements in the early Universe, and conclude that the most likely scenario to contribute to the site of elements below the third r-process peak is the magnetorotational r-process scenario, and that of the third peak is the common envelope jets supernova (CEJSN) r-process scenario. The collapsar and CEJSN r-process scenario might also contribute to the lighter r-process elements, and the binary neutron star (NS-NS) merger r-process scenario might contribute to the third r-process peak. The magnetar, the wind from the newly born NS, and the accretion-induced collapse of a white dwarf r-process scenarios fall short in explaining observations. They might exist, but cannot be major contributors to the r-process in the early Universe. To constrain r-process scenarios in the early Universe, I require that they explain the large scatter in the r-process abundances of very metal-poor stars, account for the correlation between light r-process nucleosynthesis and iron production, and the lack of correlation between the third peak r-process production and iron production, as inferred from very metal-poor stars. I discuss the diversity of the CEJSN r-process scenario and encourage extending its exploration.
I Introduction
The rapid neutron-capture process (r-process), in which intermediate-mass elements capture tens to about 200 neutrons, occurs in dynamically evolving, extremely neutron-rich sites, leading to the nucleosynthesis of the heaviest elements (e.g., Thielemann & Cowan 2026, for a recent review). There are several r-process scenarios: (1) the merger of two neutron stars, the NS-NS merger scenario (e.g., Goriely et al. 2011; Wanajo et al. 2014; Beniamini et al. 2016a, b; Beniamini & Piran 2024; Ji et al. 2016; Metzger 2017; Banerjee et al. 2020; Dvorkin et al. 2021; van de Voort et al. 2022; Maoz & Nakar 2025; Qiumu et al. 2025; Rastinejad et al. 2025); (2) the magnetorotational supernova scenario, i.e., core-collapse supernovae (CCSNe) that are powered by a fixed-axis pair of jets (e.g., Winteler et al. 2012; Nishimura et al. 2015, 2017; Halevi & Mösta 2018; Reichert et al. 2021, 2023; Yong et al. 2021; Liu et al. 2025); (3) the NS winds scenario, i.e., a wind from the newly born NS in a CCSN (e.g., Prasanna et al. 2024), possibly limited to the first r-process peak, i.e., the weak r-process (e.g., Wang & Burrows 2023, 2024), but with the third peak in the presence of magnetars (Prasanna et al., 2025); (4) the common envelope jets supernova (CEJSN) r-process scenario (e.g., Papish et al. 2015; Grichener & Soker 2019a; Grichener et al. 2022; Grichener & Soker 2022; Grichener 2023, 2025; Soker 2025); (5) collapsars, i.e., core-collapse supernova that forms a black hole at the center (e.g., Siegel et al. 2019, 2022; Brauer et al. 2021; Issa et al. 2025; Leicester et al. 2025; Mumpower et al. 2025), which probably requires significant magnetic fields (Just et al. 2022); (6) magnetar giant flares (e.g., Cehula et al. 2024; Negro et al. 2025; Patel et al. 2025a, b); (7) accretion induced collapse (AIC) of an ONeMg white dwarf (e.g., Batziou et al. 2025; Cheong et al. 2025; Combi et al. 2025; in this study, I include the merger of two white dwarfs, i.e., merger induced collapse, under AIC).
As of today, observations directly and clearly support the NS-NS merger scenario, indicating the presence of very heavy radioactive elements (e.g., Chornock et al. 2017; Kasen et al. 2017; Pian et al. 2017; Watson et al. 2019; Kasliwal et al. 2022; Watson et al. 2019; Domoto et al. 2022; Levan et al. 2024), and possibly in one magnetar giant flare (Patel et al., 2025a). However, several papers suggest that the NS-NS merger scenario may not be the sole r-process site (e.g., Waxman et al. 2018; Côté et al. 2019; Ji et al. 2019; Kobayashi et al. 2020, 2023; Holmbeck & Andrews 2024; Thielemann & Cowan 2026 for a review). Particularly, studies claimed that the lanthanide fraction (relative to total r-process yield) of kilonova GW170817/AT 2017gfo is much below the observed value, and can be more than an order of magnitude below the observed values: (Waxman et al., 2018), (Ji et al., 2019), and (Gillanders et al., 2025); most metal-poor stars have , as I present in Figure 1 adapted from Jin & Soker (2024). Many papers discuss the need for two or more r-process sites to explain the Milky Way r-process abundance (e.g., Wehmeyer et al. 2015; Grichener & Soker 2019b; Haynes & Kobayashi 2019; Molero et al. 2021; Tsujimoto 2021; Farouqi et al. 2022; Naidu et al. 2022; Yamazaki et al. 2022; Storm et al. 2025). Rastinejad et al. (2025) who study the NS-NS merger scenario, find that there is room for more r-process sites in addition to the NS-NS merger.
Some recent studies of r-process abundance in metal-poor stars claim that at least two r-process sites should have operated in the young Galaxy and Universe (e.g., Farouqi et al. 2025; Kuske et al. 2025; Saraf et al. 2025b). Other recent studies examine the scatter in r-process abundance among different metal-poor stars (e.g., Bandyopadhyay et al. 2024; Hansen et al. 2024; Xylakis-Dornbusch et al. 2024; Griffith et al. 2025; Saraf et al. 2025a). I summarize the relevant new findings of these studies (Section II) and compare the different theoretical r-process scenarios with these new observations (Section III).
The r-process scenarios are a hot topic, with numerous recent papers addressing this topic (e.g., Hayes et al. 2023; Placco et al. 2023; Anoardo et al. 2025; Bishop et al. 2025; Hansen et al. 2025; Hirai et al. 2025; Kobayashi 2025; Maoz & Nakar 2025; Molero et al. 2025; Palla et al. 2025; Santarelli et al. 2025; Storm et al. 2025; Tian et al. 2025; Xie et al. 2024, 2025). The goal of this paper is to examine actively studied r-process sites in light of recent studies of r-process nucleosynthesis in metal-poor stars, particularly in the early Universe. Namely, I address one question: Which r-process scenarios can contribute to the r-process in extremely metal-poor stars? Although this study presents no new calculations, it is the only study to systematically compare all actively studied r-process scenarios, omitting none; this allows me to conclude which scenarios are likely to contribute to r-processes in the early Universe. Therefore, it offers a timely summary of new results from 2025, and a valuable comparison of theoretical scenarios with observed r-process element properties in the early Universe. The conclusions, summarized in Section IV, have significant implications for future theoretical studies of r-process scenarios.
II R-process in metal-poor stars
To achieve my goal of listing the possible scenarios that most strongly contribute to r-process nucleosynthesis in the early Universe, I focus on two observational results from the literature, particularly those published in 2025-2026.
II.1 Large abundance scatter
From very early times of having [Fe/H], the abundance of r-process elements, like Eu, shows very large scatter, as I reproduce in Figure 2 based on Kobayashi et al. (2020). This shows that there is no averaging over many r-process nucleosynthesis events at those early times. Therefore, each event must produce a relatively large mass of r-process elements. From [Fe/H], the scatter decreases with increasing metallicity (e.g., Ou et al. 2025), indicating the averaging over more and more r-process events; above , iron production in type Ia supernovae also reduces the scatter. Henderson et al. (2025) studied 89 stars in the globular cluster M15 and found high Ba, La, and Eu dispersions in the first generation of stars. They conclude, under the assumption that the r-process events that caused the abundance dispersions were born with the first population of stars in M15, that the r-process site must have a short delay time.
These and other studies (Section I) show that all stars contain iron, which at these early times is produced by CCSNe. Iron had time to well mix with the ISM that forms these old stars. The presence of iron and the large scatter in r-process elements indicate that r-process sites are very rare relative to CCSNe. On the other hand, the presence of r-process elements indicates that the delay between CCSNe and r-process events is very short.
Racca et al. (2025) analyzed the abundance of 10 metal-poor stars and found small abundance dispersions. They concluded that this suggests a high degree of uniformity in r-process yields across diverse astrophysical environments. These stars also show large scatter in [Eu/Fe].
Kobayashi et al. (2020) argue that of that are magnetorotational supernovae that produce r-process elements can explain the r-process they deduce for the Galaxy.
Ji et al. (2023) argued that a single r-process event enriched the ultrafaint dwarf galaxy Reticulum II, and that this single event produced an r-process element mass of . More generally, they argued that the early r-process site(s) are prompt and have high yield.
The well-established conclusions from the above-cited papers and many more are as follows: (1) R-process events in the early Universe were very rare; hence, (2) they must have had a massive r-process yield per system, , and possibly a few times larger. (3) The r-process events must have a short delay from the first CCSNe.
II.2 Two early r-process sites
Farouqi et al. (2022) argue that at least three r-process nucleosynthesis sites were active in the early Galaxy: two produced Fe and the majority of the lighter r-process elements. In contrast, the third produced strong r-processed (third peak) elements. They attributed the first two to two distinct types of CCSNe, but not to standard CCSNe, and the third to NS-NS mergers, some of which rapidly produce a black hole, and some do not.
Farouqi et al. (2025) find that the Eu and Fe are co-produced at the early Universe, while the third r-process peak elements are not co-produced with Fe; the third r-process peak and Eu production are decoupled (e.g., Alencastro Puls et al. 2025). Saraf et al. (2025b) claimed the necessity of at least two r-process sites. They found that the NS-NS merger scenario cannot account for the abundance of stars with [Sr/Eu], which also tend to have a low value of [Eu/Fe]. However, they also find a continuous distribution of metal-poor stars in the plane of [Sr/Eu] versus [Eu/Fe]. Their finding is compatible with the claim of Farouqi et al. (2025) that Eu and Fe are coproduced.
I adopt these recent claims for two early r-process site types: one for light r-process elements near europium and lighter, which is correlated with iron production (I term it “Eu”+Fe), and the other for the third-peak r-process nucleosynthesis.
III Scenarios comparison
In Table 1, I compare the seven scenarios listed in Section I with respect to their ability to explain the observations related to the early r-process in the Universe described in Section II. I consider two types of r-process sites (Section II.2): one that produces the light elements in the early universe, with some correlation with iron production (I term it “Eu”+Fe), and the other that forms the third r-process peak. I also require that the scenarios operate with a short delay after the first CCSNe and account for the large scatter in r-process elements (Section II.1). A ‘’ symbol indicates that the scenario is not a main contributor to the nucleosynthesis site in the early universe. It does not imply that it does not occur. I cannot rule out that it occurs, but only that it is not the main contributer. Three of the scenarios predict that the r-process event occurs long after the CCSN; hence, in these scenarios, r-process elements are generally not correlated with iron production in the early Universe (before type Ia supernovae occur). These are indicated by ‘No’ in the third column. The other four occur alongside the CCSN explosion (or starting within seconds after the explosion).
| Scenario | RP yield | Fe | R-process: “Eu”+Fe | R-process: third peak |
| NS-NS merger | [Ra18] | No | () No Fe expected | () Ultrafast mergers[BP24]; unclear if sufficient number of systems |
| Magnetorotational | [Mo18] | Yes | () Possible | () Correlated with Fe |
| NS or magnetar wind | [Pr25] | Yes | () Too little mass to explain scatter | () Too little mass to explain scatter; () Correlated with Fe |
| CEJSN r-process | [G9S5] | No | In rare cases of a very massive oxygen core | () Possible: like a large ratio of Ir/Eu predicted[JS24] |
| Collapsar | [Si19] | Yes | () Only very strict conditions might yield sufficient mass[Is25] | () Correlated with Fe |
| Magnetar flares | [Pa25] | Yes | () Too little mass to explain scatter | () Too little mass to explain scatter |
| AIC | [Co25] | No | () Too long delay ; () No Fe expected | () Too long delay |
Notes: Comparing the different r-process scenarios during the early Universe. Additional references with further information on the properties of the different scenarios are provided in Section I. A ‘’ symbol indicates that the scenario is not the main contributor to the nucleosynthesis site in the early universe. It does not imply that it does not occur. I cannot rule out that it occurs, but only that it does not contribute substantially. I left one empty row to allow for the possibility that I unintentionally omitted a scenario. I will add it in accordance with the readers’ comments.
Columns: (1) The r-process scenario; (2) The mass of the r-process elements that one system produces; (3) The correlation with iron: A ‘Yes’ indicates that the scenario occurs during or immediately after a CCSN where iron is produced. (4) My estimate of whether a scenario can be a major contributor to the nucleosynthesis of light r-process elements (below the third peak), such as Eu, that are observed to correlate with iron in the early Universe (Section II.2); (5) My estimate of whether a scenario can be a major contributor to the third peak in the early Universe.
Abbreviations: AIC: accretion-induced collapse; CEJSN: common envelope jets supernova; NS: neutron star; RP: r-process.
Acknowledgments
I thank Aldana Grichener for enlightening conversations and suggestions. I thank the Charles Wolfson Academic Chair at the Technion for the support.